# Text2PDE: Latent Diffusion Models for Accessible Physics Simulation

### v1

@@ 5.2 Diffusion Models for PDEs
While video diffusion models can capture an understanding of physics, PDE solutions need to be more specialized and accurate than videos. In particular, neural PDE solvers have made advances in PDE operator approximation (Kovachki et al., 2021), adapting to irregular meshes or boundary conditions (Li et al., 2023c; Brandstetter et al., 2023), and accurately simulating complex, diverse physics (Alkin et al., 2024; Hao et al., 2024). Previous works on adapting diffusion models for PDEs have worked with generating individual timesteps and without the use of a latent space (Kohl et al., 2024; Lienen et al., 2024; Huang et al., 2024; Yang & Sommer, 2023). To improve temporal resolution and consistency, the use of a temporal interpolator has also been investigated (Cachay et al., 2023). Although not directly used to generate PDE solutions, diffusion has also been investigated to refine PDE solutions (Lippe et al., 2023; Serrano et al., 2024). Lastly, previous works have also considered using diffusion to accomplish alternate objectives, such as unconditional generation, upsampling PDE solutions, in-painting partial PDE solutions, or recovering solutions from sparse observations (Du et al., 2024; Shu et al., 2023)
@@ 6 Conclusion
We have demonstrated that latent diffusion models represent a powerful class of physics simulators. Specifically, we have developed a framework for encoding and decoding arbitrary PDE data into a latent space to generate unstructured physics solutions. Furthermore, we have shown that generating full spatio-temporal solutions can mitigate error accumulation to improve accuracy. Lastly, we introduce conditioning on physics data as well as language, which is a novel modality that can be more compact and universal. This creation of a text2PDE model can potentially improve the accessibility and interpretability of PDE solvers. Lastly, the scaling behavior of the proposed model is promising, opening opportunities for future work in latent diffusion or text2PDE modeling.
@@
A limitation of diffusion models is the increased time and computation required for inference. For transparency, we compare the inference times of different models on the Cylinder Flow and Smoke Buoyancy problems in Appendix G. Although out of scope for this work, there exists substantial literature on accelerating sample generation, such as distillation techniques (Salimans & Ho, 2022; Luhman & Luhman, 2021; Meng et al., 2023), faster neural ODE solvers (Song et al., 2022; 2021; Zhang & Chen, 2023; Lu et al., 2022), and consistency models (Song et al., 2023), which could be a direction for future work. An additional limitation is that the LLM captioning can sometimes hallucinate, which degrades the quality of the smoke buoyancy captions and the resulting text2PDE model. Lastly, diffusing an entire rollout at once fixes the temporal resolution of the generated solution, whereas autoregressive models can generate an arbitrarily long sequence. Similar to prior video diffusion works, this can be addressed by training a separate temporal upsampler or using the proposed model in an autoregressive fashion, albeit with a large temporal window.
@@ Acknowledgments
This research used resources of the National Energy Research Scientific Computing Center (NERSC), a Department of Energy Office of Science User Facility using a Generative AI for Science NERSC award DDR-ERCAP-m4732 for 2024.
@@ Appendix A Additional Results
Figure 5: Additional examples of text-conditioned generation of cylinder flow examples. The true solution is shown on top, followed by the sampled solution on the bottom. The model is only given a text description of the initial frame. The decoder still requires the mesh to properly collocate the solution.
@@ Appendix B Autoencoder Details
(c) Architecture Ablation for Cylinder Flow
Table 3: Validation losses between reconstructed and true samples with various autoencoder architectures.
B.1 CNN Backbone Details
We present additional results investigating the use of adversarial training and perceptual loss metrics (LPIPS) on autoencoder performance. An important consideration is that current pretrained models needed for LPIPS operate only on regularly spaced data; therefore, to maintain a fair comparison, the cylinder flow data are resampled onto a 128x32 grid to be able to evaluate an LPIPS loss. Since the benefits of using these additional methods are not clear, they are omitted from autoencoder training in the main results; however, this could be a direction for future work to improve the accuracy of LDMs. Results for the cylinder flow and smoke buoyancy problems are given in Tables 3(a) and 3(b).
Adversarial Training
A common technique to improve the reconstruction accuracy of an autoencoder is to introduce a discriminator to assess whether the reconstructed input is fake or real (Esser et al., 2021). Since our proposed mesh encoder is agnostic to the input discretization, it can also be used as a discriminator by encoding a given PDE sample, flattening the latent space, and averaging the logits. In practice, the autoencoder is trained in two stages: the first stage minimizes the reconstruction error and maximizes the probability of fooling the discriminator and the second stage optimizes the discriminator on a real and reconstructed sample.
On the cylinder flow dataset, the use of a discriminator improves autoencoder performance, however, for the smoke buoyancy dataset it does not. We hypothesize that more complex PDEs are more challenging to learn if they are fake or real, since deviations from real samples could be reasonably interpreted as eddies or high-frequency features in more turbulent regimes.
Perceptual Loss
$\lambda_{KL}$ KL Loss L1 Loss
5e-6 1.8e4 0.378
2e-7 4.8e4 0.239
Table 4: KL Weight ($\lambda_{KL}$) Ablation for Smoke Buoyancy
KL Penalty
We evaluate the effect of different KL penalties on the variance of the autoencoder latent space as well as the prediction L1 loss after downstream training of a latent diffusion model in Table 4. We find that larger KL penalties lead to lower KL losses, indicating that the latent space is closer to a Gaussian. However, this reduces the reconstruction accuracy of the autoencoder, which degrades the performance of the latent diffusion model. Alternatively, lower KL penalties result in higher-variance latent spaces, but this is outweighed by the better reconstruction capabilities.
@@ B.2 GNN/Neural Field Comparisons
We consider using GNN- and neural field-based approaches to encode and decode mesh data. The results for the proposed architectures on the cylinder flow problem are given in Table 3(c). In general, the GNN and neural field methods do not seem to perform well. We hypothesize that this could because compressing large spatio-temporal samples is very challenging without proper inductive biases. Specifically, GNNs and attention-based neural field architectures are very general in their pooling and unpooling operations, which makes it challenging to extract compressed features and reconstruct dense outputs.
@@ C.1 Denoising Ablations
Following Nichol & Dhariwal (2021) we parameterize the reverse process variance as $\Sigma_{\theta}(z_{t},t)=\exp(v\log\beta_{t}+(1-v)\log\tilde{\beta}_{t})$, where $v$ is a learned vector that modulates the transition between $\beta_{t}$ and $\tilde{\beta}_{t}$. Furthermore, the loss is modified to minimize the log-likelihood. We find that learning $\Sigma_{\theta}$ does not significantly effect the reverse process.
@@ D.1 Cylinder Flow Captioning
For the cylinder flow problem, we extract the cylinder radius, position, inlet velocity, and Reynolds number from the first frame. Additionally, we determine the flow regime based on the Reynolds number: Re<200 = laminar, Re<350 = transition, Re>350 = turbulent. Additionally, since the cylinder radius is much smaller than the domain, its dimensions are given in centimeters, while the position is given in meters. These values are inserted into the prompt below:
@@
Since the text prompt is procedurally generated, we ablate the full prompt against two baselines. The first baseline considers generating a text prompt solely with the extracted values, without any of the context or surrounding text, and uses the same RoBERTa model to encode the condition (Text-Ablate). The second baseline embeds the extracted values, normalized between -1 and 1, into a vector and uses an MLP to encode the condition (Vector-Ablate). We train these two baselines as well as a model using the full prompt and evaluate the L1 loss on the validation set in Table 7. We find that the extra context is beneficial for language models in understanding the meaning of different extracted values. Furthermore, using a language model can be a more expressive way of encoding low-dimensional physical values into a meaningful condition.
@@ D.2 Smoke Buoyancy Captioning
The captioning procedure for the smoke buoyancy dataset is more challenging, since the initial density and velocity fields do not contain features that are easily extracted and described by text. To address this, we prompt an LLM (Claude 3.5 Sonnet) to process an image of the initial density field and describe its trends. Furthermore, to improve the quality of generated captions, we use a Canny edge detector to outline shapes in the original image and provide this segmented image in the LLM prompt as well. The full prompt is given below.
@@
D.3 Classifier-Free Guidance
Classifier-free guidance can be a method to modulate the strength of the conditioning information. This is implemented by using conditional dropout during training and combining conditional and unconditional noise estimates during inference:
$\tilde{\epsilon}_{\theta}(\mathbf{z}_{n},n,\mathbf{c})=(1+w)\epsilon_{\theta}(\mathbf{z}_{n},n,\mathbf{c})-w\epsilon_{\theta}(\mathbf{z}_{n},n)$ (8)
The weight $w$ modulates the tradeoff between sample quality and diversity, with larger values increasing the guidance strength and decreasing sample diversity. Following Lin et al. (2024), we also rescale the conditional portion of the sample generation process with the recommended value of $\phi=0.7$. Empirically, we find that only at very large or small weights does the generated sample change appreciably. We plot the trends in Figure 7 for first-frame conditioning. We find that at large values of $w$, much of the generated solution is saturated and many features are washed out. At low values of $w$, many features are not present and the generated solution displays a lack of phenomena. We hypothesize that classifier-free guidance is not necessary for physics applications where the primary concern is the numerical accuracy of solutions, rather than diversity or visual quality. Indeed, the validation L1 loss of models trained without classifier-free guidance ($w=1$) is often better than using other values of $w$. The original VLB objective seeks to push generated samples closer to the training distribution; we hypothesize that modifying the inference with alternate sampling approaches deviates from this objective and produces less accurate results.

### v2

@@ 5.2 Diffusion Models for PDEs
Previous works on adapting diffusion models for PDEs have worked with generating individual timesteps and without the use of a latent space (Kohl et al., 2024; Lienen et al., 2024; Huang et al., 2024; Yang & Sommer, 2023). To improve temporal resolution and consistency, the use of a temporal interpolator has also been investigated (Cachay et al., 2023). Although not directly used to generate PDE solutions, diffusion has also been investigated to refine PDE solutions (Lippe et al., 2023; Serrano et al., 2024). Previous work has also considered using diffusion to accomplish alternate objectives, such as unconditional generation, upsampling PDE solutions, in-painting partial PDE solutions, or recovering solutions from sparse observations (Du et al., 2024; Shu et al., 2023)
@@ 6 Conclusion
We have demonstrated that latent diffusion models represent a powerful class of physics simulators. Specifically, we have developed a framework for encoding and decoding arbitrary PDE data into a latent space to generate unstructured physics solutions. Furthermore, we have shown that generating full spatio-temporal solutions can mitigate error accumulation to improve accuracy. Lastly, we introduce conditioning on physics data as well as language, which is a novel modality that can be more compact, interpretable, and accessible.
@@
A limitation of diffusion models is the increased time and computation required for inference. We seek to address this using a DDIM sampler, and show that by using fewer denoising steps, diffusion models can be faster than conventional, deterministic neural solvers while maintaining accuracy in Appendix C.3. For transparency, we also compare the inference times of using DDPM sampling on the Cylinder Flow and Smoke Buoyancy problems in Appendix G. Future work could expand on this to further accelerate sampling and improve sample quality, such as distillation techniques (Salimans & Ho, 2022; Luhman & Luhman, 2021; Meng et al., 2023) and consistency models (Song et al., 2023). An additional limitation is that the LLM captioning can sometimes hallucinate, which degrades the quality of the smoke buoyancy captions and the resulting text2PDE model. Lastly, diffusing an entire rollout at once fixes the temporal resolution of the generated solution, whereas autoregressive models can generate an arbitrarily long sequence. We present preliminary work to address this limitation by using the proposed LDM model autoregressively, to alleviate the need to scale the model when predicting longer time horizons, shown in Appendix C.4.
@@ Acknowledgments
This work was funded by Fluor Marine Propulsion, LLC under purchase order number 142474. Additionally, this research used resources of the National Energy Research Scientific Computing Center (NERSC), a Department of Energy Office of Science User Facility using a Generative AI for Science NERSC award DDR-ERCAP-m4732 for 2024.
[…]
@@ Appendix A Additional Results
Figure 4: Additional examples of text-conditioned generation of 3D turbulence samples with the velocity magnitude rendered. The true solution is shown on top, followed by the sampled solution on the bottom. While the generated solutions smoothen high-frequency features, the solutions remain stable and are broadly accurate. Rendered with vAPE4D (Koehler et al., 2024).
@@
Figure 6: Additional examples of text-conditioned generation of cylinder flow examples. The true solution is shown on top, followed by the sampled solution on the bottom.
@@ Appendix B Autoencoder Details
$\lambda_{KL}$ KL Loss L1 Loss
5e-6 1.8e4 0.378
2e-7 4.8e4 0.239
(c) KL Weight ($\lambda_{KL}$) Ablation for Smoke Buoyancy
Table 4: Validation losses between reconstructed and true samples with various CNN backbone parameters.
B.1 Ablation Studies
We present additional results investigating the use of adversarial training and perceptual loss metrics (LPIPS) on autoencoder performance. An important consideration is that current pretrained models needed for LPIPS operate only on regularly spaced data; therefore, to maintain a fair comparison, the cylinder flow data are resampled onto a 128x32 grid to be able to evaluate an LPIPS loss. Since the benefits of using these additional methods are not clear, they are omitted from autoencoder training in the main results; however, this could be a direction for future work to improve the accuracy of LDMs. Results for the cylinder flow and smoke buoyancy problems are given in Tables 4(a) and 4(b).
Adversarial Training
A common technique to improve the reconstruction accuracy of an autoencoder is to introduce a discriminator to assess whether the reconstructed input is fake or real (Esser et al., 2021). Since our proposed mesh encoder is agnostic to the input discretization, it can also be used as a discriminator by encoding a given PDE sample, flattening the latent space, and averaging the logits. In practice, the autoencoder is trained in two stages: the first stage minimizes the reconstruction error and maximizes the probability of an erroneous discriminator classification, and the second stage optimizes the discriminator when given a real and reconstructed sample.
On the cylinder flow dataset, the use of a discriminator improves autoencoder performance, however, for the smoke buoyancy dataset it does not. We hypothesize that more complex PDEs are more challenging to learn if they are fake or real, since deviations from real samples could be reasonably interpreted as eddies or high-frequency features in more turbulent regimes. We anticipate this could be improved with additional tuning, but due to its difficulty we omit the use of adversarial training for the main results, and indeed lower losses can also be obtained but simply scaling the autoencoder and training it for longer.
Perceptual Loss
KL Penalty
We evaluate the effect of different KL penalties on the variance of the autoencoder latent space as well as the prediction L1 loss after downstream training of a latent diffusion model in Table 4(c). We find that larger KL penalties lead to lower KL losses, indicating that the latent space is closer to a Gaussian. However, this reduces the reconstruction accuracy of the autoencoder, which degrades the performance of the latent diffusion model. Alternatively, lower KL penalties result in higher-variance latent spaces, but this is outweighed by the better reconstruction capabilities.
Ball Radius L1 Loss
0.045 .0138
0.033 .0141
0.020 .0148
(a) Ball Radius Ablation for Cylinder Flow
@@ B.1 Ablation Studies
(b) Architecture Ablation for Cylinder Flow
Table 5: Validation losses between reconstructed and true samples with various mesh learning parameters.
Ball Radius
We consider training the mesh encoder/decoder with different ball radii to measure the effect of the size of receptive field of the learnable aggregation and interpolation. Using the same hyperparameters for the CNN backbone, we train the mesh autoencoder at different ball radii on the cylinder flow benchmark and report the validation L1 reconstruction losses in Table 5(a). We find that increasing the radius improves the reconstruction accuracy, likely since there are more points within the ball used to construct the latent grid, as well as more latent grid points used to query output points. However, this comes at an additional computational cost due to the need to integrate over more points when approximating the kernel integral.
@@ B.2 GNN/Neural Field Comparisons
We consider using GNN- and neural field-based approaches to encode and decode mesh data. The results for the proposed architectures on the cylinder flow problem are given in Table 5(b). In general, the GNN and neural field methods do not seem to perform well. We hypothesize that this could because compressing large spatio-temporal samples is very challenging without proper inductive biases. Specifically, GNNs and attention-based neural field architectures are very general in their pooling and unpooling operations, which makes it challenging to extract compressed features and reconstruct dense outputs.
@@
Figure 7: Kernel Aggregation. Neighbors to a latent grid points are used to compute an operator mapping between the input mesh and latent grid. This can also be extended to arbitrary dimensions, such as including a time dimension or extra space dimension. Additionally, this can be reversed to compute an operator mapping between the decoded latent grid and output mesh.
B.3 Kernel Aggregation/Interpolation
[…]
Figure 8: Discretization Invariance. Given a sampled latent solution from the conditional LDM, it can be decoded onto an arbitrary set of query points.
@@ C.1 Denoising Ablations
Following Nichol & Dhariwal (2021) we investigate parameterizing the reverse process variance as $\Sigma_{\theta}(z_{t},t)=\exp(v\log\beta_{t}+(1-v)\log\tilde{\beta}_{t})$, where $v$ is a learned vector that modulates the transition between $\beta_{t}$ and $\tilde{\beta}_{t}$. Furthermore, the loss is modified to minimize the log-likelihood. We find that learning $\Sigma_{\theta}$ does not significantly effect the reverse process.
@@
Model S Inf. Time Rel. L2 Loss
GINO - 0.112 0.2445
MGN - 1.683 0.2617
OFormer - 1.336 0.3386
LDMS-FF 10 0.4572 0.1595
20 0.6119 0.1583
50 1.0844 0.1579
100 1.8716 0.1595
1000 15.753 0.1522
LDMM-FF 10 0.5672 0.1313
20 0.8654 0.1267
50 1.7188 0.1306
100 3.1561 0.1281
1000 27.895 0.1309
LDMS-Text 10 0.4711 0.1941
20 0.6405 0.1980
50 1.1522 0.1949
100 1.9936 0.1903
1000 16.251 0.1796
LDMM-Text 10 0.5770 0.1474
20 0.8836 0.1536
50 1.7635 0.1513
100 3.2456 0.1495
1000 29.096 0.1476
Table 7: Cylinder Flow Sampling. Comparison of inference times (s) and relative L2 errors of different baselines and LDM models under varying numbers of denoising steps S using DDIM or DDPM sampling.
@@ C.2 Diffusion Backbone Architectures
C.3 DDIM Sampling
We evaluate the use of DDIM to accelerate sampling after training an LDM with the denoising objective. We implement a DDIM sampler (Song et al., 2022) and evaluate the inference time and relative L2 loss of sampled solutions on the Cylinder Flow problem at different denoising steps S. These results are presented in Table 7 on a single NVIDIA A100 GPU. We find that DDIM can recover most of the solution even at a small number of denoising steps, and in certain cases produce higher quality samples. Through the use of DDIM, in combination with using a latent space and spatio-temporal sampling, diffusion models can produce solutions faster than conventional, deterministic neural solvers in addition to being more accurate. We hypothesize that DDIM is effective due to fundamental differences in the modeled distributions within the image and PDE domains.
We consider the underlying conditional distribution $p_{\theta}(\mathbf{u}|\mathbf{c})$ that is approximated by conditional diffusion. Within image domains, this distribution has many modes in order to cover a diverse set of sampled images for a given prompt, however, in PDE settings this distribution ideally only has a single mode since when conditioned on a past timestep, there should only be a single sampled solution. Therefore, in PDE settings, we can expect Langevin sampling to converge quickly, since the distribution that is sampled from has only a single mode. Furthermore, since score estimates are still trained using the original denoising objective, they are still accurate at the initial sample $\mathbf{x}_{0}$; therefore, although Langevin sampling can start far from $p_{\theta}(\mathbf{u}|\mathbf{c})$, the score estimates are still accurate enough to quickly approach the desired mode.
One method of empirically evaluating this could be to calculate the Jacobian of the estimated score function $\nabla_{x}s_{\theta}(x)$ for diffusion models trained in PDE and image settings. The Jacobian of the score function would correspond to the Hessian of the estimated underlying potential function $\nabla^{2}_{x}F_{\theta}(x)$, which is related to the mean mixing time of the Langevin sampler. The mean mixing time is correlated with the inverse of the Hessian determinant (Bovier et al., 2004), which could measure the complexity of the estimated distribution $p_{\theta}(x)\propto F_{\theta}(x)$ and its effect on sampling times.
Model Rel. L2 Loss
GINO Unstable
MGN 0.171
OFormer 0.088
LDMAR-FF 0.068
LDMAR-Text 0.085
Table 9: Left: Autoregressive LDMM results on Cylinder Flow, using identical hyperparameters to Table 1. Right: Framework for autoregressively generating text-conditioned PDE samples.
C.4 Autoregressive LDMs
Although sampling complete PDE trajectories can result in lower error accumulation, this fixes the temporal horizon and would require scaling the model in order to achieve longer rollouts. To address this, we present results on using the proposed LDM framework autoregressively, albeit with a large prediction window. For LDM models conditioned on physical field values, this can be easily done by passing the last frame of the predicted solution back to the LDM model. However, for text-conditioned LDM models, the prompt is only relevant for the first prediction. To address this, we investigate a conditioning strategy in which the model can dynamically switch between text- and physics-based conditioning during inference by padding the unused modality with zeros, shown in Table 9. From a user perspective, querying the model with a text prompt remains the same; however, the model internally uses different conditioning modalities at different steps.
To validate this, we train the LDM model on the Cylinder Flow problem to predict 100 timesteps, in four windows of 25 timesteps each. We also compare this approach with neural solver baselines re-trained with a longer prediction horizon. To maintain a comparison to the original experiment, we keep the hyperparameters and model sizes the same and present the results in Table 9 for the medium-sized LDM model. Predictably, errors grow across all models in the longer time horizon setting, with GINO becoming unstable due to error propagation. Notably, the text-conditioned LDM model incurs a performance loss presumably due to switching modalities during inference.
@@ D.1 Cylinder Flow Captioning
For the cylinder flow problem, we extract the cylinder radius, position, inlet velocity, and Reynolds number from the first frame. Since the cylinder radius is much smaller than the domain, it is given in centimeters while the position is given in meters. Additionally, we determine the flow regime based on the Reynolds number: Re<200 = laminar, Re<350 = transition, Re>350 = turbulent. One limitation is that this is overly simplistic; flow regimes are generally problem-specific and can vary even with the same Reynolds number. In addition, there are varying standards to characterize these flow regimes from Reynolds numbers (Blevins, 1990; Zdravkovich, 1990; Stringer et al., 2014a). However, given the lack of resources to manually label this dataset, we adopt this simple strategy since these semantic differences do not greatly affect performance. Values are inserted in the prompt:
@@
Since the text prompt is procedurally generated, we ablate the full prompt against two baselines. The first baseline considers generating a text prompt solely with the extracted values, without any of the surrounding text, and uses the same RoBERTa model to encode the condition (Text-Ablate). The second baseline embeds the extracted values, normalized between -1 and 1, into a vector and uses an MLP to encode the condition (Vector-Ablate). We train these two baselines and a model using the full prompt and report the validation L1 loss in Table 10. We find that the extra context is beneficial for language models in understanding the meaning of different extracted values. Furthermore, using a language model can be a more expressive way of encoding low-dimensional physical values into a meaningful condition.
@@ D.2 Smoke Buoyancy Captioning
The captioning procedure for the smoke buoyancy dataset is more challenging, since the initial density and velocity fields do not contain features that are easily extracted and described by text. To address this, we prompt an LLM (Claude 3.5 Sonnet) to describe an image of the initial density field. To improve the quality of generated captions, we also use a Canny edge detector to outline shapes in the image and provide this segmented image in the LLM prompt. The prompt is given below.
@@
D.3 3D Turbulence Captioning
Due to the small number of samples, we manually caption this dataset. To maintain consistency across human-generated captions, we follow a template that asks for a pre-specified set of features. In particular, we look to describe qualitative features, the position, and the size of the shape, as well as how many shapes are present. An example template is provided below; however, the templates can vary based on the quantity and composition of the shapes:
"Air flows over an obstacle resembling a {qualitative_label}. The {shape} is {shape_label} and is located at {position_label}. The flow is turbulent."
@@
D.4 Classifier-Free Guidance
Classifier-free guidance can be a method to modulate the strength of the conditioning information. This is implemented by using conditional dropout during training and combining conditional and unconditional noise estimates during inference: $\tilde{\epsilon}_{\theta}(\mathbf{z}_{n},n,\mathbf{c})=(1+w)\epsilon_{\theta}(\mathbf{z}_{n},n,\mathbf{c})-w\epsilon_{\theta}(\mathbf{z}_{n},n)$. The weight $w$ modulates the tradeoff between sample quality and diversity, with larger values increasing the guidance strength and decreasing sample diversity. Following Lin et al. (2024), we also rescale the conditional portion of the sample generation process with the recommended value of $\phi=0.7$. Empirically, we find that only at very large or small weights does the generated sample change appreciably. We plot the trends in Figure 10 for first-frame conditioning. We find that at large values of $w$, much of the generated solution is saturated and many features are washed out. At low values of $w$, many features are not present and the generated solution displays a lack of phenomena. We hypothesize that classifier-free guidance is not necessary for physics applications where the primary concern is the numerical accuracy of solutions, rather than diversity or visual quality and indeed the loss of models trained without classifier-free guidance ($w=1$) is often better than using other values of $w$.